OpenStudy (anonymous):

a car is travelling at 60 km/h. How far will it take in a) 3 hours? b) 30 minutes? c) 2 minutes?

OpenStudy (anonymous):

Since D=ST, a) 180 Km b) 30 Km c) 2 Km

OpenStudy (anonymous):

THANXX :)

OpenStudy (anonymous):

s = 60 * 3 = 180 km s = 60 *1/2 = 30km s = 60 * 2/60 = 2km